Grade 11Mechanics

a ball is moving eastward with speed 35m/s an acc 10m/s2 is acting on it find the distance covered in 4th sec

anil kumar
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
From the formulae:
x = u + a*(2n-1)/2 we get,
x = 35 + 10*(2*4 – 1)/2 => 35 + 35 = 70 meter.
This is the distance the ball would cover in east direction.